# Exchange control in a MOS double quantum dot made using a 300 mm wafer process

### v2

@@ Acknowledgements
In radio-frequency reflectometry, the measured signal is directly proportional to the change in capacitance of the system, this can be described mathematically as;
$\Delta C_{j}=\frac{dQ_{T}}{dV_{j}}$ (17)
Here $Q_{T}$ denote the total charge of the system and $V_{j}$ represents the $j^{th}$ gate. We take inspiration from [53] and rewrite the change in capacitance as measured from gate $V_{j}$ to be;
$\Delta C_{j}=\sum_{i=1}^{M}C_{i,j}=e\sum_{i=1}^{M}\alpha_{i,j}\frac{d\langle n_{i}\rangle}{dV_{j}}$ (18)
In this context, $C_{i,j}$ represents the capacitance felt from dot $i$ by gate $j$. The average occupancy of dot $i$ is denoted $\langle n_{i}\rangle$. The lever arm matrix is defined as $\hat{\alpha}$ as $\hat{\alpha}=\mathbf{C_{cc}^{-1}}\mathbf{C_{cv}}$, which links gate-induced potential changes to the charge states of the quantum dots.
To calculate $\langle n_{i}\rangle$, we iterate through each Fock state in $\mathcal{F}$ and compute its corresponding probability. Subsequently, a weighted sum of the occupation numbers of each Fock state at position $i$ is performed, as expressed below:
@@
where $P_{k}$ represents the probability of the quantum dot array being in the Fock state $\Lambda_{k}$.
We assume a Boltzmann distribution and write the probability accordingly;
@@
where $Z$ represents the partition function, and $\epsilon_{k}$ represents the energy required for quantum dot array to be in the Fock state $\Lambda_{k}$. Substituting this into equation 19, we obtain;
@@
This term can be further substituted into equation 18, to obtain;
@@
We use the above formalism to simulate the charge stability diagrams in the $V_{\text{G}_{1}}-V_{\text{G}_{2}}$ gate-voltage space. This can be seen in Fig. 5f-i, where different voltage configurations were applied to the multi-electron dot via $V_{\text{G}_{\text{S}}}$.

### v3

@@ Methods
In radio-frequency reflectometry, the measured signal is directly proportional to the change in capacitance of the system, this can be described mathematically as
$\Delta C_{j}=\frac{dQ_{T}}{dV_{j}},$ (24)
here $Q_{T}$ denotes the total charge of the system and $V_{j}$ represents the $j^{th}$ gate. We take inspiration from [74] and rewrite the change in capacitance as measured from gate $V_{j}$ to be
$\Delta C_{j}=\sum_{i=1}^{M}C_{i,j}=e\sum_{i=1}^{M}\alpha_{i,j}\frac{d\langle n_{i}\rangle}{dV_{j}}.$ (25)
In this context, $C_{i,j}$ represents the capacitance felt from dot $i$ by gate $j$. The average occupancy of dot $i$ is denoted $\langle n_{i}\rangle$. The lever arm matrix is defined as $\alpha=\mathbf{C_{cc}^{-1}}\mathbf{C_{cv}}$, which links gate-induced potential changes to the charge states of the quantum dots. To calculate $\langle n_{i}\rangle$, we iterate through each Fock state in $\mathcal{F}$ and compute its corresponding probability. Subsequently, a weighted sum of the occupation numbers of each Fock state at position $i$ is performed, expressed as
@@
where $P_{k}$ represents the probability of the quantum dot array being in the Fock state $\Lambda_{k}$. We assume a Boltzmann distribution and write the probability accordingly;
@@
where $Z$ represents the partition function, and $\epsilon_{k}$ represents the energy required for quantum dot array to be in the Fock state $\Lambda_{k}$. Substituting this into eq. (26), we obtain;
@@
This term can be further substituted into eq. (25), to obtain;
@@
We use the above formalism to simulate the charge stability diagrams in the $V_{\text{G}_{1}}-V_{\text{G}_{2}}$ gate-voltage space. This can be seen in Supplementary Fig. S1(f-j), where different voltage configurations were applied to the multi-electron dot via $V_{\text{G}_{\text{S}}}$.
Data Availability
Acknowledgements
We acknowledge helpful conversations with H. Jnane, A. Seigel, S.C. Benjamin, A.J. Fisher and G. Burkard at Quantum Motion. We also acknowledge technical support from G. Antilen Jacob at the London Centre for Nanotechnology. This work received support from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 951852 (Quantum Large Scale Integration in Silicon); from the Engineering and Physical Sciences Research Council (EPSRC) under grant Nos. (EP/S021582/1), (EP/L015978/1), (EP/T001062/1) and (EP/L015242/1); and from Innovate UK under grant Nos. (43942) and (10015036). T.M. acknowledges support from the Winton Programme for the Physics of Sustainability. M.F.G.Z. acknowledges support from the UKRI Future Leaders Fellowship (MR/V023284/1).
Author contributions
J.F.C.W. conducted the experiments and analysed the results presented in this work with input from R.C.C.L., M.A.F., M.F.G.Z. and J.J.L.M.; J.F.C.W., M.A.F. and R.C.C.L. conducted preliminary experiments; T.M. and G.A.O developed the quantum dot rf simulator in Supplementary Section 4 informed by electrostatic modelling from J.W.; T. M. performed and wrote about the simulations under the supervision of D. F. W. and M. F. G. Z.; S.M.P. designed the device under the supervision of M.A.F., M.F.G.Z. and J.J.L.M.; J.J. and S.K. fabricated the device under the supervision of B.G.; F.E.v.H. characterised the superconducting resonator under the supervision of M.F.G.Z.; N.J. maintained the experimental setup; J.F.C.W., M.F.G.Z. and J.J.L.M. wrote the manuscript with input from R.C.C.L., M.A.F., N.J. and S.M.P.; M.F.G.Z. and J.J.L.M. conceived and oversaw the experiment.
Competing Interests
Quantum Motion Technologies Limited has filed a patent application related to this work: EP23206749.6 (Europe) by J.F.C.W., R.C.C.L., M.A.F. and M.F.G.Z.